how many sandwiches can you make with 5 meats 4 breads and 4 cheeses
we think the answer is 20 or 80 but book doesn't provide example for a 3 tier choice or 2 so were a lil confused

Found 2 solutions by rfer, jim_thompson5910:
Answer by rfer(16322)

your answer is 80.

Answer by jim_thompson5910(35256)

The answer is 5*4*4 = 20*4 = 80